# Divide-and-Conquer Dynamics in AI-Driven Disempowerment

### v1

@@ 2.1 Models for the battle-win probability $p_{i}$
We first consider the simple case in which Movement 1’s battle-win probability, $p_{i}$, does not depend on the state $i$, i.e., $p_{i}=p$ for all $i\in I$. Then, the win probability of Movement 1, $q$, can be explicitly calculated [23, Section 12.2] as
@@ 3 Game theory: Cooperate or not?
We assume that there is an extra up-front cost $c>0$ that each coalition member needs to pay if and only if it agrees to this new larger coalition. We also assume that if Movement 1 wins, each of its members gain a periodic benefit $b$ for all future time periods $\ell$, which (using an exponential discounting of $e^{-r\ell}$, as is standard in economics) produces a total benefit of $\sum_{\ell=1}^{\infty}be^{-r\ell}=b/(e^{r}-1)$. We allow for $b$ to be either positive, indicating that the member benefits from their movement prevailing; or negative, indicating that the member is in fact harmed by their movement prevailing.
@@
We assume that each member is rational, in that when they choose their action in the set $A=\{S,G\}$, where $S$ denotes ‘voting for the status quo’ and $G$ denotes ‘voting for greater unity,’ they are assumed to maximize their expected payoff. This motivates the definition of a (pure) Nash equilibrium, a vector of the players’ actions
@@ 4.3 Divide-and-conquer dynamics in military conflicts
Napoleon’s unusually high rate of military victory is often attributed to his ability “to overcome a superior enemy by dividing their forces.” [19, p. 3] This is especially well-demonstrated by his first Italian campaign during the War of the First Coalition. In it, Napoleon adeptly employed speed, maneuverability, and central positioning of his 38,000 troops to divide the defending forces, comprised of 52,000 Piedmontese and Austrian forces. [25] By defeating the former into surrendering and subsequently vanquishing the latter, Napoleon was able to achieve a complete victory over a numerically superior adversary.
@@
Divide-and-conquer dynamics were of pivotal significance in the First Balkan War. This is well-explained by the following preamble quote from Edward Erickson [21, p. xvii]:
@@ 5.2 Copyright lawsuits
“I am looking for lawyer with experience in filing DCMA claims, specifically with regards to tools that ‘circumvent technological protection measures […] that control access to copyrighted works.’…
@@ 5.3 Contract negotiations and strikes for AI-ban clauses
This is the last time any of these unions will have leverage. We see how much has happened in the last, what has it been, two months? In three years, we won’t have this kind of leverage. So if we want to get anything for the next three years, we’ve got to go all in.
History suggests that WGA’s contract win will help set the tone for SAG-AFTRA’s contractual negotiations,[61] an illustrative example of how different stakeholders of the pro-human movement can cooperate synergistically.
@@ 5.5 Political advocacy for AI regulation
One can argue that all the previous battlefields ultimately lead to this arguably most important battlefield. Stakeholders can advocate for the adoption of laws and regulations that presciently guide AI development, and impose robust restrictions on its application to use cases that threaten indefinite human economic uselessness. Stakeholders may, for example, advocate for the establishment of legal, regulatory, governmental, and inter-governmental infrastructures pertaining to the use of advanced AI in society.[47] Advocates of AI deregulation, in contrast, would advocate for the continued absence of such laws and regulations. Who prevails in this debate will be largely determined by the degree to which the voters, lobbying actors, and politicians comprising the pro-human movement are united or divided—for instance, in their efforts to advocate for supportive policies—relative to those comprising the AI replacement movement. If the former triumphs over the latter, temporary wins—such as the WGA’s contractual protections against AI encroachment—can be elevated to permanent wins at the level of government policy.

### v2

@@ 2.1 Models for the battle-win probability $p_{i}$
We first consider the simple case in which Movement 1’s battle-win probability, $p_{i}$, does not depend on the state $i$, i.e., $p_{i}=p$ for all $i\in I$. Then, the win probability of Movement 1, $q$, can be explicitly calculated [24, Section 12.2] as
@@ 3 Game theory: Cooperate or not?
We assume that there is an extra up-front cost $c>0$ that each coalition member needs to pay if and only if they agree on this new larger coalition. We also assume that if Movement 1 wins, each of its members gains a periodic benefit $b$ for all future time periods $\ell$, which (using an exponential discounting of $e^{-r\ell}$, as is standard in economics) produces a total benefit of $\sum_{\ell=1}^{\infty}be^{-r\ell}=b/(e^{r}-1)$. We allow for $b$ to be either positive, indicating that the member benefits from their movement prevailing; or negative, indicating that the member is in fact harmed by their movement prevailing.
@@
We assume that each member is rational, in that when they choose their action in the set $A=\{S,G\}$, where $S$ denotes ‘voting for the status quo’ and $G$ denotes ‘voting for greater unity,’ they are assumed to maximize their expected payoff. This motivates the definition of a (pure) Nash equilibrium, [43] a vector of the players’ actions
@@ 4.3 Divide-and-conquer dynamics in military conflicts
Napoleon’s unusually high rate of military victory is often attributed to his ability “to overcome a superior enemy by dividing their forces.” [20, p. 3] This is especially well-demonstrated by his first Italian campaign during the War of the First Coalition. In it, Napoleon adeptly employed speed, maneuverability, and central positioning of his 38,000 troops to divide the defending forces, comprised of 52,000 Piedmontese and Austrian forces. [26] By defeating the former into surrendering and subsequently vanquishing the latter, Napoleon was able to achieve a complete victory over a numerically superior adversary.
@@
Divide-and-conquer dynamics were of pivotal significance in the First Balkan War. This is well-explained by the following preamble quote from Edward Erickson [22, p. xvii]:
@@ 5.2 Copyright lawsuits
“I am looking for [a] lawyer with experience in filing DCMA claims, specifically with regards to tools that ‘circumvent technological protection measures […] that control access to copyrighted works.’…
@@ 5.3 Contract negotiations and strikes for AI-ban clauses
“This is the last time any of these unions will have leverage. We see how much has happened in the last, what has it been, two months? In three years, we won’t have this kind of leverage. So if we want to get anything for the next three years, we’ve got to go all in.”
WGA’s contract win also helped set the tone for SAG-AFTRA’s contract win,[2] an illustrative example of how different stakeholders of the pro-human movement can cooperate synergistically.
@@ 5.5 Political advocacy for AI regulation
One can argue that all the previous battlefields ultimately lead to this arguably most important battlefield. Stakeholders can advocate for the adoption of laws and regulations that presciently guide AI development, and impose robust restrictions on its application to use cases that threaten indefinite human economic uselessness. Stakeholders may, for example, advocate for the establishment of legal, regulatory, governmental, and inter-governmental infrastructures pertaining to the use of advanced AI in society.[49] Advocates of AI deregulation, in contrast, would advocate for the continued absence of such laws and regulations. Who prevails in this debate will be largely determined by the degree to which the voters, lobbying actors, and politicians comprising the pro-human movement are united or divided—for instance, in their efforts to advocate for supportive policies—relative to those comprising the AI replacement movement. If the former triumphs over the latter, temporary wins—such as the WGA’s and SAG-AFTRA’s contractual protections against AI encroachment—can be elevated to permanent wins at the level of government policy.